# Sequential model confidence sets

### v1

@@ 3.2 SMCSs for the strongly superior models
Alternatively, one could adjust the $m\cdot(m-1)$ e-processes $(E_{ij,t})_{t\in\mathbb{N}}$ directly to obtain FWV e-processes $(E^{\star}_{ij,t})_{t\in\mathbb{N}}$ for the hypotheses $\mathcal{H}^{\textrm{s}}_{ij}$ and exclude model $i$ from the SMCS at $t\in\mathbb{N}$ if $E^{\star}_{ij,t}>1/\alpha$ for some $j\neq i$. However, the corresponding tests are less powerful, since for any $i\in\mathcal{M}_{0}$ and $j\neq i$,
$E^{\star}_{ij,t}$ $=\min\left\{\frac{1}{\lvert I\rvert}\sum_{(k,l)\in I}E_{kl,t}\;\Big{|}\;I\subseteq\{1,\dots,m\}^{2},(i,j)\in I\right\}$
Remark 2.
Apart from the FWER, the false discovery rate (FDR) is arguably the most commonly used criterion in multiple testing (Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001). Recently, Wang and Ramdas (2022) studied FDR control for e-values, which turn out to be particularly suitable for this problem due to their behaviour under taking expectations. In Appendix D, we demonstrate how one could alternatively construct SMCSs with a time-uniform control of the FDR, which is generally less restrictive than FWER.
3.3 SMCSs for the uniformly weakly superior objects
We construct SMCSs for the sequence of uniformly weakly superior objects $(\mathcal{M}^{\textrm{uw},\star}_{t})_{t\in\mathbb{N}}$ defined at (2) analogously as in the previous section. For simplicity, we assume first that for any $i,j\in\mathcal{M}$, there exists a constant $c_{ij}>0$ with $\lvert d_{ij,t}\rvert\leq c_{ij}/2$ for all $t\in\mathbb{N}$. In Section 3.5, we explain how we can weaken this assumption to conditionally bounded loss differences.
For $i,j\in\mathcal{M}$, consider the weak hypotheses
$\mathcal{H}^{\textrm{uw}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\Delta_{ij,t}\leq 0,\forall t\in\mathbb{N}\}$
and the weak intersection hypotheses
$\mathcal{H}^{\textrm{uw}}_{i\cdot}=\bigcap_{j\in\mathcal{M}_{0}}\mathcal{H}^{\textrm{uw}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\Delta_{ij,t}\leq 0,\forall t\in\mathbb{N},\forall j\in\mathcal{M}_{0}\}.$
Choe and Ramdas (2023) show that, for any $0\leq\lambda_{ij}<1/c_{ij}$,
$M_{ij,t}=\exp\left\{\lambda_{ij}t(\hat{\Delta}_{ij,t}-\Delta_{ij,t})-\psi_{E,c_{ij}}(\lambda_{ij})V_{ij,t}\right\},\quad t\in\mathbb{N}$
is a test supermartingale under any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, where
$\psi_{E,c}(\lambda)=(-\log(1-c\lambda)-c\lambda)/c^{2}$ (11)
for $c>0$ and $0\leq\lambda<1/c$ and
$V_{ij,t}=\sum_{r=1}^{t}(d_{ij,r}-\gamma_{ij,r})^{2}$ (12)
for some predictable sequence $(\gamma_{ij,t})_{t\in\mathbb{N}}$ with $\lvert\gamma_{ij,t}\lvert\leq c_{ij}/2$. Importantly, for any $\mathbb{Q}\in\mathcal{H}^{\textrm{uw}}_{ij}$, $M_{ij,t}$ upper bounds
$E_{ij,t}=\exp\left\{\lambda_{ij}t\hat{\Delta}_{ij,t}-\psi_{E,c_{ij}}(\lambda_{ij})V_{ij,t}\right\}$ (13)
for all $t\in\mathbb{N}$. That is, $(E_{ij,t})_{t\in\mathbb{N}}$ is an e-process for $\mathcal{H}^{\textrm{uw}}_{ij}$ and hence also for $\mathcal{H}^{\textrm{uw}}_{i\cdot}$. Analogously as before, we adjust the averages $E_{i\cdot,t}=1/(m-1)\sum_{j\neq i}E_{ij,t}$, $i=1,\dots,m$, symmetrically by the closure principle to construct SMCSs for $(\mathcal{M}^{\textrm{uw},\star}_{t})_{t\in\mathbb{N}}$ by the FWV e-processes $(E_{1\cdot,t}^{\star})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t}^{\star})_{t\in\mathbb{N}}$ as in (10), compare Appendix C.
3.4 SMCSs for the weakly superior objects
Our SMCS construction for the weakly superior objects defined at (3) builds on time-uniform confidence regions. For any $i,j\in\mathcal{M}_{0}$, we again assume that there exists $c_{ij}>0$ with $\lvert d_{ij,t}\rvert\leq c_{ij}/2$ for all $t\in\mathbb{N}$ and discuss a possible weakening of this assumption in Section 3.5. For $t\in\mathbb{N}$ and $0\leq\lambda_{ij}\leq 1/c_{ij}$, let
$M_{ij,t}(x)=\exp\left\{\lambda_{ij}t\hat{\Delta}_{ij}-\lambda_{ij}tx-\psi_{E,c_{ij}}(\lambda_{ij})V_{ij,t}\right\},\quad x\in\mathbb{R},t\in\mathbb{N},$
where $\psi_{E,c_{ij}}$ and $V_{ij,t}$ are given in (11) and (12), respectively. The following lemma shows the one-to-one correspondence between confidence sequences and sequential tests, and is found e.g. in Howard et al. (2021, Lemma 1) or Waudby-Smith and Ramdas (2024, Theorem 1).
Lemma 3.2.
For any $i,j\in\mathcal{M}_{0}$ and $\alpha\in(0,1)$, the sequence
$C_{t}=\left\{x\in[-c_{ij}/2,c_{ij}/2]\mid M_{ij,t}(x)\leq 1/\alpha\right\},\quad t\in\mathbb{N},$
is a $(1-\alpha)$-confidence sequence for $(\Delta_{ij,t})_{t\in\mathbb{N}}$.
The claim follows directly by Ville’s inequality observing that $(M_{ij,t}(\Delta_{ij,t}))_{t\in\mathbb{N}}$ is a nonnegative supermartingale for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$. Note that $M_{ij,t}(x)\leq 1/\alpha$ if and only if $x\geq\hat{\Delta}_{ij,t}-(\psi_{E,c_{ij}}(\lambda_{ij})V_{ij,t}+\log(1/\alpha))/(\lambda_{ij}t)$. That is, Lemma 3.2 actually yields time-uniform lower confidence bounds for $(\Delta_{ij,t})_{t\in\mathbb{N}}$. Next, we show how to combine these lower confidence bounds simultaneously for all models $i\in\mathcal{M}_{0}$, and how we can use them to construct SMCSs for $(\mathcal{M}^{\textrm{w},\star}_{t})_{t\in\mathbb{N}}$.
We let $\mathbb{R}_{0}^{m\times m}$ be the family of all $m\times m$-matrices with diagonal entries equal zero and let
$\boldsymbol{c}=\begin{pmatrix}0&c_{12}&\cdots&c_{1m}\\ c_{21}&0&\cdots&c_{2m}\\ \vdots&\vdots&\ddots&\vdots\\ c_{m1}&c_{m2}&\dots&0\\ \end{pmatrix}\in\mathbb{R}_{0}^{m\times m},\quad\boldsymbol{\Delta}_{t}=\begin{pmatrix}0&\Delta_{12,t}&\cdots&\Delta_{1m,t}\\ \Delta_{21,t}&0&\cdots&\Delta_{2m,t}\\ \vdots&\vdots&\ddots&\vdots\\ \Delta_{m1,t}&\Delta_{m2,t}&\cdots&0\\ \end{pmatrix}\in\mathbb{R}_{0}^{m\times m},\quad t\in\mathbb{N}.$
For $\boldsymbol{A},\boldsymbol{B},\boldsymbol{C}\in\mathbb{R}_{0}^{m\times m}$ and $C\in\mathbb{R}$, we understand $\boldsymbol{A}\leq\boldsymbol{B}$ and $C\cdot\boldsymbol{A}$ componentwise, and write $\boldsymbol{B}\leq\boldsymbol{A}\leq\boldsymbol{C}$ equivalently as $\boldsymbol{A}\in[\boldsymbol{B},\boldsymbol{C}]$. For $\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}$, we let
Lemma 3.3.
For any $\alpha\in(0,1)$, the sequence
$C_{t,1-\alpha}=\left\{\boldsymbol{X}\in[-\boldsymbol{c}/2,\boldsymbol{c}/2]\mid M_{t}(\boldsymbol{X})\leq 1/\alpha\right\}$
is a time-uniform $(1-\alpha)$-confidence sequence for the sequence $(\boldsymbol{\Delta}_{t})_{t\in\mathbb{N}}$.
Proof.
For any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, we have
$\mathbb{Q}(\exists t\in\mathbb{N}:\boldsymbol{\Delta}_{t}\notin C_{t,1-\alpha})$ $=\mathbb{Q}(\exists t\in\mathbb{N}:M_{t}(\boldsymbol{\Delta}_{t})>1/\alpha)\leq\alpha,$
by Ville’s inequality since $(M_{t}(\boldsymbol{\Delta}_{t}))_{t\in\mathbb{N}}$ is a test supermartingale under any $\mathbb{Q}\in\mathfrak{B}(\Omega)$. ∎
Next, we define
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid C_{t,1-\alpha}\cap\left(\mathbb{R}_{0}^{m\times m}\right)_{ij,-}\neq\emptyset\textrm{ for all }j\neq i\right\},\quad t\in\mathbb{N},$ (14)
where $\left(\mathbb{R}_{0}^{m\times m}\right)_{ij,-}=\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid X_{ij}\leq 0\}$. That is, for each model $i\in\mathcal{M}_{0}$, we check whether $\Delta_{ij,t}$ may potentially be negative for all $j\neq i$ according to the confidence region $C_{t,1-\alpha}$. If there exists $j\neq i$ such that we may reject that $\Delta_{ij,t}\leq 0$ with high probability, then we exclude $i$ from the SMCS.
Lemma 3.4.
For any $\alpha\in(0,1)$, the sequence (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ defined at (14) is a a sequence of model confidence sets at level $\alpha$ for the sequence of weakly superior objects $(\mathcal{M}^{\textrm{w},\star}_{t})_{t\in\mathbb{N}}$.
Proof.
By Lemma 3.3, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, we have
$\mathbb{Q}(\exists t\in\mathbb{N}:\mathcal{M}^{\textrm{w},\star}_{t}\not\subseteq\widehat{\mathcal{M}}_{t})$
$=$ $\mathbb{Q}(\exists t\in\mathbb{N}:i\in\mathcal{M}^{\textrm{w},\star}_{t}\textrm{ and }i\not\in\widehat{\mathcal{M}}_{t}\textrm{ for some }i\in\mathcal{M}_{0})$
$=$ $\mathbb{Q}(\exists t\in\mathbb{N},\exists i\in\mathcal{M}_{0}:\Delta_{ij,t}\leq 0\textrm{ for all }j\neq i\textrm{ and }C_{t}\cap\left(\mathbb{R}^{m\times m}\right)_{i\ell,-}=\emptyset\textrm{ for some }\ell\neq i)$
$\leq$ $\mathbb{Q}(\exists t\in\mathbb{N}:\boldsymbol{\Delta}_{t}\notin C_{t})\leq\alpha.$
The following Lemma, which is proved in Appendix E, characterizes the shape of the confidence regions $(C_{t,1-\alpha})_{t\in\mathbb{N}}$ and may be used to compute the sequence (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ more efficiently.
Lemma 3.5.
For any $\alpha\in(0,1)$ and $t\in\mathbb{N}$, $C_{t,1-\alpha}$ is a convex upper set. That is, for $\boldsymbol{A},\boldsymbol{B}\in C_{t,1-\alpha}$, it follows that $\lambda\boldsymbol{A}+(1-\lambda)\boldsymbol{B}\in C_{t,1-\alpha}$ for all $\lambda\in[0,1]$, and $\boldsymbol{A}\in C_{t,1-\alpha}$ and $\boldsymbol{A}\leq\boldsymbol{C}$ implies that $\boldsymbol{C}\in C_{t,1-\alpha}$ for any $\boldsymbol{C}\leq\boldsymbol{c}/2$.
By Lemma 3.5, we may express ^$\widehat{\mathcal{M}}_{t}$ equivalently as
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid C_{t,1-\alpha}\cap[\boldsymbol{0},\boldsymbol{c}/2]_{ij,0}\neq\emptyset\textrm{ for all }j\neq i\right\},\quad t\in\mathbb{N},$
where $\boldsymbol{0}=(0)_{i,j=1}^{m}\in\mathbb{R}_{0}^{m\times m}$ and $[\boldsymbol{0},\boldsymbol{c}/2]_{ij,0}=\{\boldsymbol{X}\in[\boldsymbol{0},\boldsymbol{c}/2]\mid X_{ij}=0\}$. That is, we exclude model $i\in\mathcal{M}_{0}$ from the SMCS at $t\in\mathbb{N}$, if there exists $j\neq i$ such that the confidence region $C_{t,1-\alpha}$ has an empty intersection with the hyperplane $\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid X_{ij}=0\}$.
3.5 Predictable bounds and betting schemes
The methods provided in Sections 3.3 and 3.4 require that the score differences are uniformly bounded over time, which is clearly more restrictive than the assumption of conditionally bounded score differences imposed in Section 3.2. However, one can always transform conditionally bounded score differences into uniformly bounded differences thereby modifying the underlying loss and the corresponding superior objects.
Assume that $\lvert d_{ij,t}\rvert\leq c_{ij,t}/2$ for all $t\in\mathbb{N}$ for some predictable $(c_{ij,t})_{t\in\mathbb{N}}$. Then, the transformed loss differences $\tilde{d}_{ij,t}=d_{ij,t}/c_{ij,t},t\in\mathbb{N}$ are uniformly bounded with constant $c=1$ that is $\lvert\tilde{d}_{ij,t}\rvert\leq 1/2$ for all $t\in\mathbb{N}$. However, if we use the transformed loss differences for the methods of the previous sections, then we target (possibly) different sequences of superior objects
~$\tilde{\mathcal{M}}^{\textrm{uw},\star}_{t}=\left\{i\in\mathcal{M}_{0}\mid\tilde{\Delta}_{ij,r}\leq 0,\forall r\leq t,\forall j\in\mathcal{M}_{0}\right\},\quad\tilde{\mathcal{M}}^{\textrm{w},\star}_{t}=\left\{i\in\mathcal{M}_{0}\mid\tilde{\Delta}_{ij,t}\leq 0,\forall j\in\mathcal{M}_{0}\right\},$
for
$\tilde{\Delta}_{ij,t}=\frac{1}{t}\sum_{r=1}^{t}\tilde{\mu}_{ij,r}=\frac{1}{t}\sum_{r=1}^{t}\mathbb{E}(\tilde{d}_{ij,r}\mid\mathcal{F}_{r-1})=\frac{1}{t}\sum_{r=1}^{t}\frac{\mu_{ij,r}}{c_{ij,r}}.$
In Section 5, we use the given transformation to convert conditionally bounded CRPS differences of wind gust forecasts into uniformly bounded loss differences. Importantly, for any proper scoring rule (consistent scoring function) ${\rm S}$, the scaled function $\tilde{{\rm S}}={\rm S}/c$ is proper (consistent) as well for any $c>0$, an observation which justifies the transformation from a theoretical perspective. Additionally, with the given scaling, all observations have the same maximal impact on the ranking, an effect which seems appealing for many applications.
Remark 3.
We assume conditionally bounded loss differences in order to work with the empirical Bernstein e-processes at (13) provided by Choe and Ramdas (2023). Clearly, for many important loss functions this assumption is not fulfilled; see, e.g., Example 2.2 where the log-likelihood differences are typically unbounded. However, if the loss differences emerge from (or may be bounded in the tails by) a parametric family of distributions, as it is for example the case in Example 2.2, there are other well-studied e-processes available; see, e.g., Howard et al. (2020), Pérez-Ortiz et al. (2023) or Hao et al. (2024) amongst others. In such situations, our methods directly extend with a suitable choice of e-processes at (13).
We conclude with a comment on the choice of the parameter $\lambda$. For the uniformly weak hypothesis and the weak hypothesis, we assume a universal bound $c>0$ on the loss differences, and $\lambda$ may be any fixed value in the interval $(0,c^{-1})$. If there is no good reason to do otherwise, we suggest to choose $\lambda=(2c)^{-1}$ by default. In betting language, this corresponds to bet half of the accumulated evidence at each time step (Shafer, 2021). For the strong hypothesis, we should choose $\lambda$ predictably and in a smart way in order to increase power. Clearly, there are many possible betting schemes. In Section 5.1, we propose a particular betting scheme which makes use of all previous observations and the assumption that the relative performance of two models does not change too quickly over time. Another option would be the method of mixtures (Robbins, 1970), where we integrate over all possible values of $\lambda$ over a particular probability distribution instead of choosing one specific value. Mixtures of e-processes are again e-processes and may be expressed in closed form for some distributions (Howard et al., 2021, Appendix A.3). The method of mixtures is one of the most widely-studied techniques for constructing uniform bounds which are shown to be unimprovable under certain conditions (Robbins and Siegmund, 1970, Howard et al., 2021). However, often it is analytically opaque and computationally tedious, and does not lead to closed-form confidence sequences. In accordance with Waudby-Smith and Ramdas (2024), for the strong hypothesis, our experience shows that choosing $\lambda$ predictably in a smart ways is generally more promising than the methods of mixtures. Also for the weaker hypotheses, we could not improve the power of our methods significantly by using mixtures instead of choosing a fixed $\lambda$ value. For this reason, and to make the computations more tractable, we suggest to work with the given e-processes directly rather than with the more complicated mixture e-processes also for both weak hypotheses.
@@ 4.1 Simulation 1
We consider the setup given in Example 2.3. That is, we consider $n=1000$ realizations from a sequence $Y_{t}\sim\mathcal{N}(Y_{t-1},1)$ with $Y_{0}=0$, and assess $m=49$ forecasters which issue sequentially probabilistic predictions with respect to the CRPS. We write $Y_{t}=Y_{t-1}+Z_{t}$, for an i.i.d.  standard normally distributed sequence $(Z_{t})_{t\in\mathbb{N}}$, and use the formulas provided by Grimit et al. (2006) to derive
$L_{i,t}$ $=\textrm{CRPS}(f_{i,t},Y_{t})$
$=\sqrt{1+\delta_{i}}\left(\frac{Z_{t}-\varepsilon_{i}}{\sqrt{1+\delta_{i}}}\left(2\Phi\left(\frac{Z_{t}-\varepsilon_{i}}{\sqrt{1+\delta_{i}}}\right)-1\right)+2\varphi\left(\frac{Z_{t}-\varepsilon_{i}}{\sqrt{1+\delta_{i}}}\right)-\frac{1}{\sqrt{\pi}}\right),$ (15)
where $\Phi$ and $\varphi$ denote the cdf and the density function of the standard normal distribution, respectively. Let $(\mathcal{F}_{t})_{t\in\mathbb{N}}$ be the canonical filtration, that is $\mathcal{F}_{t}$ is generated by all observations $Y_{r}$ and forecasts $f_{i,r}$, $i=1,\dots,m$, $r\leq t$. For any $t\in\mathbb{N}$, we have $\mathbb{E}(d_{ij,t}\mid\mathcal{F}_{t-1})=\mathbb{E}(d_{ij,t})=\mathbb{E}(d_{ij,1})$ by independence. Since the CRPS is a proper scoring rule, we conclude
$\mathcal{M}^{\textrm{s},\star}_{t}=\mathcal{M}^{\star}=\left\{i\in\mathcal{M}_{0}\mid\mathbb{E}(d_{ij,1})\leq 0,\forall j\in\mathcal{M}_{0}\right\}=\{i_{0}\}$
for the forecaster $i_{0}$ with $\varepsilon_{i_{0}}=\delta_{i_{0}}=0$.
We construct SMCSs for $(\mathcal{M}^{\textrm{s},\star}_{t})_{t\leq n}$ as proposed in Section 3.2 with confidence level $\alpha=0.1$. We let $\lambda_{ij,t}=(2c_{ij})^{-1}$ for the universal bound $c_{ij}>0$ on the score differences $(d_{ij,t})$ given in Appendix B. Figure 1 shows the average size of the SMCS over time for $N=100$ realizations of the simulation. Importantly, we have a 100% coverage rate of the best model $i_{0}$, that is, in all runs of the simulation, the SMCS never wrongly excludes $i_{0}$.
Figure 1: The average number of models in the SMCS in simulations 1 and 2 over $N=100$ repetitions. In simulation 1, the size of the SMCS tends to the actual number of superior objects as the sample size increases, whereas there are still approximately $10$ models remaining in the SMCS for the uniformly weakly superior objects after $1000$ observations. For both simulations, the SMCS never wrongly excludes the best model $i_{0}$.

### v3

@@ 3.4 Predictable bounds and betting schemes
For important choices of loss functions such as the quadratic or logarithmic score, the corresponding loss differences are not conditionally bounded. However, if the loss differences emerge from (or may be bounded in the tails by) a parametric family of distributions, there are other well-studied e-processes available; see, e.g., Howard et al. (2020). In Supplement F.1, we provide a simulation example of mean forecasts with Gaussian errors. The resulting score differences are unbounded but still sub-exponential, which allows to use the e-processes of Proposition 3.3.
We conclude with a comment on the choice of the parameters $\lambda_{ij}$. For the uniformly weak and weak hypothesis (Propositions 3.3 and 3.6), we assume universal bounds $c_{ij}>0$ on the pairwise loss differences, and $\lambda_{ij}$ may be any fixed value in the interval $(0,c_{ij}^{-1})$. As a default, we suggest to choose $\lambda_{ij}=(2c_{ij})^{-1}$. In betting language, this corresponds to bet half of the accumulated evidence at each time step (Shafer, 2021). For the strong hypothesis (Proposition 3.2), we can choose $\lambda_{ij,t}$ predictably to increase power. There are many possible betting schemes. In Section 5.1, we propose a particular betting scheme which makes use of all previous observations and the assumption that the relative performance of two models does not change too quickly over time. Another option would be the method of mixtures (Robbins, 1970), where we integrate over all possible values of $\lambda_{ij,t}$ over a particular probability distribution instead of choosing one specific value. Mixtures of e-processes are again e-processes and may be expressed in closed form for some distributions (Howard et al., 2021, Appendix A.3). The method of mixtures is one of the most widely-studied techniques for constructing uniform bounds which are shown to be unimprovable under certain conditions (Robbins and Siegmund, 1970, Howard et al., 2021). However, in accordance with Waudby-Smith and Ramdas (2024), for the strong hypothesis, our experience shows that choosing $\lambda_{ij,t}$ predictably is generally more promising than the methods of mixtures. Also for the weak hypotheses, where predictable betting strategies do not apply, we could not improve the power of our methods significantly by using mixtures instead of choosing $\lambda_{ij}$ fixed. For this reason, and to make computations more tractable, we suggest to work with the given e-processes directly.
@@ 4.1 Simulation 1
We consider the setup given in Section 2.1. That is, we consider $n=1000$ realizations from a sequence $Y_{t}\sim\mathcal{N}(Y_{t-1},1)$ with $Y_{0}=0$, and assess $m=49$ forecasters which issue sequentially probabilistic predictions with respect to the CRPS. We write $Y_{t}=Y_{t-1}+Z_{t}$, for an i.i.d.  standard normally distributed sequence $(Z_{t})_{t\in\mathbb{N}}$, and use the formulas provided by Grimit et al. (2006) to derive
$L_{i,t}=\sqrt{1+\delta_{i}}\left(\frac{Z_{t}-\varepsilon_{i}}{\sqrt{1+\delta_{i}}}\left(2\Phi\left(\frac{Z_{t}-\varepsilon_{i}}{\sqrt{1+\delta_{i}}}\right)-1\right)+2\varphi\left(\frac{Z_{t}-\varepsilon_{i}}{\sqrt{1+\delta_{i}}}\right)-\frac{1}{\sqrt{\pi}}\right),$ (11)
where $\Phi$ and $\varphi$ denote the cdf and density function of the standard normal distribution, respectively. Let $(\mathcal{F}_{t})_{t\in\mathbb{N}}$ be the canonical filtration, that is, $\mathcal{F}_{t}$ is generated by $Y_{r}$ and $f_{i,r}$ for $i=1,\dots,m$, $r\leq t$. For any $t\in\mathbb{N}$, we have $\mathbb{E}(d_{ij,t}\mid\mathcal{F}_{t-1})=\mathbb{E}(d_{ij,t})=\mathbb{E}(d_{ij,1})$ by independence. Since the CRPS is a proper scoring rule, we conclude $\mathcal{M}^{\textrm{s},\star}=\left\{i\in\mathcal{M}_{0}\mid\mathbb{E}(d_{ij,1})\leq 0,\forall j\in\mathcal{M}_{0}\right\}=\{i_{0}\}$ for the forecaster $i_{0}$ with $\varepsilon_{i_{0}}=\delta_{i_{0}}=0$.
We construct SMCSs for $\mathcal{M}^{\textrm{s},\star}$ with confidence level $\alpha=0.1$. Upper bounds on the score differences $(d_{ij,t})_{t\in\mathbb{N}}$ are given in Supplement B. These do not depend on time, since biases and dispersion errors do not change, so $c_{ij,t}=c_{ij}>0$, $t\in\mathbb{N}$. Define $\lambda_{ij,t}=(2c_{ij})^{-1}$.
Figure 1 shows the average size of the SMCS over time for $N=1000$ simulations. We have a 100% coverage rate for the best model $i_{0}$, that is, in all simulation runs, the SMCS never incorrectly excludes $i_{0}$. Hence, our methods are conservative. However, using e-processes and a sequential application of the closure principle, we cannot expect to improve the power of our methods for SMCSs, since the arithmetic average, which is the e-merging function that we used, essentially dominates any other e-merging function.
Figure 1: The average number of models in the SMCS in Simulation 1 and 2. At the end of the evaluation period, the SMCSs have an average size of $8.41$ and $9.95$, respectively. For both simulations, the SMCS never wrongly excludes the best model $i_{0}$.
We ran Simulation 1 also for the original MCS by Hansen et al. (2011). Although its size is generally better than that of the SMCSs, the coverage rate drops below the nominal level $0.91-\alpha=$ 0.9 to about $0.8$. The sequential implementation of the MCS is documented in Supplement F.2, along with another simulation in a more challenging, non-Gaussian setting. There, the coverage rates of the MCS are heavily compromised, in contrast to those of the SMCS.